A temporary agency supplies temporary employees to other companies each workday. They pay an average of $10.00/hour and charge an average of $16.00/hour. Each temporary worker averages 160 hours per month on a job. For the office and its staff they have fixed operating costs of $18.00/month. In the month of February the company had a $10,542 net gain. How many employees did they have in February? (Be sure to show all of your work)
this is the problem that i am having hard time to solved.

Answer by mananth(16949)

COSTS
Let there be x workers in february
$10 per hour
160 hours per worker
so cost per month for all workers 10*160*x
=$1600x
Fixed cost $18
Total cost = 1600x+18
..
INCOME
Company charges $16/worker/ hour
160 hours = 16*160= $2560 per worker
for x workers income = 2560x
..
Net gain = $10,542
..
Income - expenditure = gain
..
2560x-(1600x+18)= 10,542
2560x-1600x-18=10542
960x=10542+18
960x= 10560
x= 10560/960
x= 11 workers